Grade 12General Physics

The work of 146 kJ is performed in order to compress one kilo mole of gas adiabatically and in this process the temperature of the gas increases by 7° C. The gas is ?

sudhanshu
12 Years agoGrade 12

1 Answer

Aatif nisar
9 Years ago
Answer is DiatomicDQ=0Du=-∆wnCvdt=146* 10^3Nfr\2*7=146*10^3F=5Hence diatomicF is degree of freedom